• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ulmer

    Log entry time 00:08:12 on November15,2002

    Entry number 90111

    keyword=Swing Shift Summary

    No beam at start of shift - MCC needs to reboot IOC.

    16:05: MCC calls. They will deliver low current and ramp up.
    16:13: MCC says Hall A ep ion chamber is causing beam trips. They will look into it.
    16:15: MCC says ep ion chamber fault now clear - beam is on the way.
    16:18: MCC says ion chamber tripped again, even before beam could be ramped up. Radcon has been called.
    16:20: MCC calls: Radcon needs to go into hall to investigate ion chamber. We will go to controlled access.
    17:08: MCC will try to deliver beam. They are now monitoring the ep ion chamber dose rate with a strip tool and will now be able to take corrective action should the dose rate increase.
    17:18: After about 15 uA, MCC calls: they want to steer beam to reduce scrape-off. So they're sending tuned beam now.
    17:25: Low current CW on the way.
    17:50: MCC: beam still tripping ep ion chamber. The position near this chamber is -3.0 mm (at 1H01) according to MCC. This seems low, but is apparently where we have been running for several days. So what's changed? In any case, they will try to move it to slightly larger position 1h01 (about -1.0mm). Now they get 100 uA for a while but they have other trips: RF and injector.
    18:30: MCC: lots of injector trips. So the problems pile up.
    19:00: MCC reports that they are ready to send beam.
    19:16: 100 uA beam. BPMB is nicely centered, but BPMAX has opposite sign from before (+0.6 mm vs. -0.4 mm). We take it and will check the data. Starting run 3172.
    20:34: Target problem - automatic control failed and temperature began to fluctuate. Ask for beam OFF. J.P. Chen (target on-call) was called.
    21:23: Manual control of target (with advice from J.P.) has managed to get stable temperature at 22.00 deg. Switched to automatic control and asked for beam back.
    See Halog 90105.
    21:35: Beam back. It took 12 minutes. Apparently, the ep ion chamber tripped when our slits were opened, even before beam was sent. Perhaps there's something wrong with the device.


    Run Events Sum-trues (for all runs up to listed run)

    3172 1.0M 15
    3173 0.3M 16
    3174 1.0M 20
    3175 1.0M 24

    Average: 0.030 trues/uA-hr (into cuts)
    Goal: 140 trues